I really couldn't find a catagory for my pre- Algebra problem sorry.
Well here is my problem x-21=-2. Thank you for your time:)
x=-2+21
x=19

Found 2 solutions by Juhi, Niku:
Answer by Juhi(44)

x - 21 = -2
x = - 2 + 21
= 19

Answer by Niku(6)

x=-2+21
x=19